The following is the information from the national income accounts for a hypothetical country: GNP Rs. 5000.00 Personal Disposable Income 4100.00 Consumption 3800.00 X-M 50.00 Govt. Budget Deficit 200.00 Calculate Gross Investment and Government Expenditure
1
Expert's answer
2016-11-05T12:12:08-0400
GNP = Y = Rs. 5000.00, Personal Disposable Income DI = 4100.00, Consumption C = 3800.00, X-M = 50.00, Govt. Budget Deficit BD = 200.00.
DI = Y - T, BD = G - T, Y = C + I + G + X - M, so: Government Expenditure G = BD + T = BD + (Y - DI) = 200 + 5000 - 4100 = 1100.
As Y = C + G + I + X - M, then Gross Investment I = Y - C - G - (X - M) = 5000 - 3800 - 1100 - 50 = 50.